Find the value of root (i) + root (-i).
I solved this but I didn't understand why the imaginary value were not considered as per the answer to the question is given.
It's urgent please help me out as soon as possible

Grade:11

1 Answers

Aditya Gupta
5 years ago
 root (i) + root (-i)= root (i) + 1/root(i) [because –i = 1/i, which implies root–i = root1/i]
=(i+1)/root(i)
we know that root(i)= ±(1+i)/root2
so, root (i) + root (-i)= ±root2